OpenStudy (anonymous):

A child leaps on a skateboard: A 35 kg girl is running at a speed of 2.8 m/s when she jumps on a stationary skateboard. If the system consisting of the girl and the skateboard begins rolling at a speed of 2.6 m/s, what is the mass of the skateboard?

OpenStudy (anonymous):

use the momentum of conservation formula

OpenStudy (anonymous):

M1V1+M2V2=M3V3 M1V1=(M1+M2)V3 IE 35*2.8=(35+x)2.6 x=2.69 kilos

OpenStudy (anonymous):

V2=0 since at the beginning the skateboard is at rest

OpenStudy (anonymous):

cheers glad to help you dont forget to fan me